OpenStudy (anonymous):

which equation is a standard form and represents a line with slope -1 through the point (-3,5)? ...-x+y=5 ...y=-x+2 ...y=-x+5 ...x+y=2

OpenStudy (aroub):

y=mx+b 5=-1(-3)+b 5=3+b b=2 y=-x+2